OpenStudy (emily_98):

Which statement describes the graph of f(x) = 4x2 + 20x + 25? The graph does not intersect the x-axis. The graph touches the x-axis at (–2.5, 0). The graph intersects the x-axis at (–0.4, 0) and (0.4, 0). The graph intersects the x-axis at (2, 0) and (5, 0).

OpenStudy (raffle_snaffle):

First graph the function in a graphing calculator.

OpenStudy (greatlife44):

Intercepts the x axis, pretty sure that's when y = 0. \[4x^{2}+20x+25 = 0 \] \[-b \pm \frac{ \sqrt{b^{2}-4ac} }{ 2a }\] a = 4, b = 20 c = 25 You could apply the quadratic formula to find where where f(x) = 0. this will give you values where y = 0. \[\sqrt{20^{2}-4(4)(25)} = 0 \] \[\frac{ -20+0 }{ 8 } = \frac{ -5 }{ 2 } = (x_{1})\]

OpenStudy (emily_98):

the second one???

OpenStudy (greatlife44):

(-2.5,0) the idea is that you apply the quadratic formula